A 9​-year bond pays interest of $ 28.60 ​semiannually, has a face value of $1,000​, and...

A

9​-year

bond pays interest of

$ 28.60

​semiannually, has a face value of

$1,000​,

and is selling for

$765.73.

What are its annual coupon rate and yield to​ maturity?

The annual coupon rate is

__​%.

​(Round to two decimal​ places).
Answer #1

1)

Annual coupon = 28.6 * 2 = 57.20

Annual coupon rate = (Annual coupon / face value) * 100

Annual coupon rate = (57.20 / 1,000) * 100

Annual coupon rate = 5.72%

2)

Number of periods = 9 * 2 = 18

Yield to maturity = 9.68%

Keys to use in a financial calculator:

2nd P/Y 2

FV 1000

PV -765.73

PMT 28.6

N 18

CPT I/Y